OpenStudy (anonymous):

What are the real or imaginary solutions of each polynomial equation? x^4-29x^2+100

OpenStudy (anonymous):

(x^2-29)x^2+100 (x-5)(x-2)(x+2)(x+5) (x^229/2)^2-441/4

OpenStudy (anonymous):

x=-5 x=-2 x=2 x=5

OpenStudy (anonymous):

x^4-29x^2+100 = 0 let u=x^2. so your equation can look like this quadratic: u^2 - 29u +100 = 0

OpenStudy (anonymous):

factoring you get (u-25)(u-4) = 0 so u = 25 or u = 4 so x^2 = 25 or x^2 = 4

OpenStudy (anonymous):

so x = 5 or -5 x = 2 or -2